OpenStudy (anonymous):

prove each identity: tan theta = 1-cos2theta / sin 2 theta

OpenStudy (anonymous):

use 1-cos 2theta=2sin^2 theta sin 2 theta=2 sin theta cos theta and solve

OpenStudy (anonymous):

im not getting it right, can you write it for me using the drawing?

OpenStudy (anonymous):

\[\tan \theta =\frac{ 1-\cos 2\theta }{ \sin 2\theta } \] \[1-\cos 2\theta=2\sin ^{2}\theta ,\sin 2\theta =2\sin \theta \cos \theta \] solve it